# Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Writing - original draft, Writing - review & editing, Visualization

### v2

@@ 3.4.1 Pre-training: Barlow Twins Slide Mixing
Initializing neural networks with pre-trained models offers the advantage of leveraging knowledge from unlabeled datasets, providing a strong starting point with weights that capture general patterns and representations. This strategy is particularly beneficial for tasks with limited labeled data [45].
@@
Slide augmentation. Standard augmentations from the PyTorch Torchvision library [46] are not directly applicable in our context, as our inputs are extracted WSI features rather than natural images. To address this limitation, we adapt four specific augmentations—random zeroing, Gaussian noise addition, random scaling, and random cropping—so they can be applied to WSI features during pre-training. As shown in Figure 3, for the $N^{\text{th}}$ WSI in the batch, the augmented features are denoted as $X_{N\text{A}}$ and $X_{N\text{B}}$. After applying augmentations to all WSIs in the batch, the features $X_{1\text{A}}$ to $X_{N\text{A}}$ are combined into $X_{\text{A}}$, and similarly, $X_{1\text{B}}$ to $X_{N\text{B}}$ are combined into $X_{\text{B}}$.
Intra-batch mixing. Inspired by prior work [42], which demonstrated the effectiveness of image mixing in MoCo V2 [30] for improving downstream image classification, we introduce an intra-batch mixing strategy to the Barlow Twins method (a non-contrastive approach) to enhance the generation of positive slide pairs during MIL aggregator pre-training. Specifically, slide mixing is performed within a single batch to produce additional inputs, denoted as $X_{A-\text{Mix}}$ and $X_{B-\text{Mix}}$, as illustrated in Figure 3.
@@
The details of the OriginalLoss and MixingLoss functions, used to compute the Barlow Twins Slide Mixing loss, are provided in Algorithm 1. The OriginalLoss function calculates the standard Barlow Twins loss, while the MixingLoss function computes the loss between $Z_{B-{\text{Mix}}}$ and $Z_{A}$ or $Z_{B-{\text{Mix}}}$ and $Z_{A-{\text{Mix}}}$, accounting for $\lambda(i)$ as defined above. Here, "Flip" refers to reversing the batch order.
source_loss $\leftarrow\textsc{OriginalLoss}(Z_{A},Z_{B})$ (9)
The total loss is defined as follows:
@@ 4 Experiments
Table 1: WSI dataset distribution during pre-training and downstream classification phases. An "x" indicates that the dataset is not used in a particular process (e.g., pre-training, training, or testing), while a "-" indicates that distributions depend on the random sampling or active learning (AL) strategy used to select a 20 WSI sample budget per iteration from the unlabeled pool.
@@ 4.1 Data acquisition and preparation
This study utilizes the Camelyon16 dataset [47] [^1] Link to Camelyon16 dataset, which contains 399 hematoxylin and eosin (H&E) stained whole slide images (WSIs) of breast cancer lymph node sections collected from two hospitals in the Netherlands: Radboud University Medical Center (RUMC) and University Medical Center Utrecht (UMCU). The dataset is divided into 270 training slides (159 normal and 111 tumor) and 129 test slides, serving as a widely used benchmark for WSI classification tasks.
For the pre-training phase, we supplement Camelyon16 with data from the Clinical Proteomic Tumor Analysis Consortium (CPTAC) Uterine Corpus Endometrial Carcinoma (UCEC) dataset [48] [^2] Link to CPTAC UCEC dataset. From this dataset, 575 H&E stained WSIs of endometrial tissue samples are selected, comprising 185 normal and 390 tumor slides. The UCEC dataset focuses on endometrial cancer, a significant gynecological malignancy, providing diverse tissue samples for robust representation learning. During pre-training, both Camelyon16 and UCEC datasets are utilized in an unsupervised manner, with no access to slide-level labels to ensure a purely self-supervised learning process. All test data are strictly excluded from the pre-training phase to prevent information leakage.
To evaluate the model’s performance in data-limited scenarios, we restrict training to 100 WSIs from the Camelyon16 dataset. Initially, 20 WSIs are randomly selected for training, followed by iterative sample selection using either random sampling or active learning acquisition functions [49, 50, 51, 52, 53]. The complete Camelyon16 test set of 129 WSIs is retained for final evaluation. A detailed summary of the WSI dataset distribution across the pre-training and downstream classification phases is provided in Table 1.
@@ 4.2 Implementation setup
To operate under resource constraints and demonstrate the efficiency of our framework, all training experiments, including pre-training and downstream classification, are conducted on a single NVIDIA GeForce RTX 2080 Ti GPU. The models are implemented and trained using the PyTorch library [46], a widely used open-source deep learning platform, with all experiments performed in Python. For tiling WSIs into distinct regions, we use the OpenSlide library [54].
@@ 4.2.1 Region extraction
WSIs are tiled into regions of size $[4096,4096]$ pixels with no overlap between adjacent regions. A tissue threshold of 0.1 is applied to exclude non-informative regions. To standardize tissue patch extraction and address magnification differences across scanners in the Camelyon16 dataset, we employ HS2P [55], based on CLAM [12], with a pixel-spacing set to 0.5 μm/px (equivalent to 20x magnification).
@@ 4.3 Model evaluation
In the baseline framework, the MIL aggregator is initialized and trained from scratch. To evaluate the effectiveness of the proposed framework, we compute the average F1 for each active learning acquisition function and random sampling (traditional fully-supervised fine-tuning) across label budgets. This comprehensive evaluation highlights the robustness and adaptability of our framework under varying dataset compositions and labeling budgets.
@@ 4.4 Impact of MIL aggregator pre-training methods
Building on the findings of [16], which employs the contrastive SimCLR method for MIL aggregator pre-training, we evaluate the non-contrastive Barlow Twins method alongside our proposed Barlow Twins Slide Mixing approach. The results, summarized in Table 2, reveal a clear difference in performance between these methods. Pre-training the MIL aggregator with SimCLR results in a 3.8% reduction in mean F1 compared to the baseline (training the aggregator from scratch), averaged across random sampling (traditional fully-supervised fine-tuning) and five active learning acquisition functions. In contrast, the Barlow Twins method achieves a 1.4% improvement over the baseline, highlighting the advantages of non-contrastive approaches for this task.
Our proposed Barlow Twins Slide Mixing method enhances the vanilla Barlow Twins approach by introducing intra-batch mixing, which generates additional positive WSI pairs during pre-training. This improvement results in a further 2% increase in mean F1 compared to the vanilla Barlow Twins method. These results underscore the effectiveness of our pre-training strategy in improving WSI classification by leveraging better initialization methods using unlabeled WSIs.
@@ 4.5 Impact of Mixup and Manifold Mixup integration
Table 3 reports the average performance across random sampling (traditional fully-supervised fine-tuning) and five active learning acquisition functions under five different WSI training label budgets. Integrating Mixup and Manifold Mixup during the MIL aggregator fine-tuning phase, following Barlow Twins Slide Mixing initialization, achieves the highest performance. Specifically, this approach delivers a substantial improvement of 1.3% compared to the configuration without Mixup and Manifold Mixup and a total improvement of 4.7% over the baseline. The performance gains are consistent across different acquisition strategies and budget constraints, demonstrating the robustness of our framework, particularly in scenarios with limited labeled WSIs. These findings emphasize the critical role of proper initialization through pre-training and the significant benefits of incorporating Mixup and Manifold Mixup during fine-tuning, ultimately contributing to enhanced WSI classification.
Table 3: Overall Results: Average performance of various acquisition functions under different budget constraints, evaluated with the F1 score. ’SM’ means Slide Mixing.
@@ 4.6.1 Effect of slide mixing loss hyperparameters
We investigate the impact of adjusting the hyperparameters $(\alpha,\beta,\gamma)$ in the Barlow Twins Slide Mixing method over 300 epochs. These hyperparameters control the relative weights of $source\_loss$, $mix\_source\_loss$, and $mix\_loss$ in Equation 10. As summarized in Table 4, the optimal performance is achieved with hyperparameter values set to $(1,0.5,0.5)$. Other configurations yield slightly lower mean F1, suggesting that the performance of Barlow Twins Slide Mixing is robust to moderate variations in loss hyperparameters. These results highlight the method’s flexibility and effectiveness across a range of parameter settings.
Table 4: Effect of adjusting hyperparameters $(\alpha,\beta,\gamma)$ in Barlow Twins Slide Mixing loss, corresponding to $source_{l}oss$, $mix_{s}ource_{l}oss$, and $mix_{l}oss$ in Equation 10, respectively, evaluated with the F1 score.
@@ 4.6.2 Effect of Mixup and Manifold Mixup locations
We analyze the impact of applying Mixup and Manifold Mixup at different locations during MIL aggregator fine-tuning. This evaluation focuses on the performance of mixing at $Loc1$, $Loc2$, and $Loc3$, as illustrated in Figure 4. According to Table 5, incorporating mixing at all three locations achieves the highest mean F1 of 74.6%, outperforming other configurations. In contrast, applying mixing only at $Loc1$ results in a slight performance decline, underscoring the importance of leveraging multiple mixing locations. Notably, placing Manifold Mixup at $Loc2$, within the Transformer Encoder layer, proves particularly advantageous, yielding a mean F1 improvement of 2.5%. These findings highlight the synergistic benefits of comprehensive mixing across layers, with the Transformer Encoder layer playing a pivotal role in driving the greatest performance gains.
@@ 5 Discussion
Most existing works train the MIL aggregator from scratch, overlooking the valuable information present in unlabeled WSIs. Our findings indicate that fine-tuning the MIL aggregator with a contrastive method, such as SimCLR, performs worse than supervised training in improving WSI classification. This trend is consistent with prior work [16]. The primary limitation of contrastive methods lies in their reliance on negative pairs alongside positive pairs. This approach is unsuitable for pre-training the MIL aggregator at the WSI level because the inherent imbalance in gigapixel WSI datasets (e.g., a significantly larger number of normal WSIs compared to tumor WSIs) often results in the inclusion of incorrect or noisy negative pairs. To address this issue, we recommend adopting non-contrastive methods such as Barlow Twins, which rely solely on positive pairs. Our experimental results demonstrate that, unlike contrastive methods that perform worse than the baseline (training the MIL aggregator from scratch), the non-contrastive approach significantly enhances WSI classification performance.
@@
The proposed Barlow Twins Slide Mixing method further builds on the vanilla Barlow Twins approach by introducing an intra-batch mixing strategy that generates additional positive slide pairs during pre-training. This strategy enriches the pre-training process by leveraging the semantic relatedness of mixed slide features, even when the mixed features originate from different slides within the same batch. This additional semantic diversity enables the MIL aggregator to learn more robust and meaningful representations.
To improve downstream performance, we integrate Mixup and Manifold Mixup, designed to handle the unique gigapixel nature and variable sizes of WSIs. Unlike their original implementations for natural images of uniform size, our adaptations interpolate features and labels across diverse WSI dimensions. These techniques serve as effective augmentation methods, generating synthetic WSIs that enhance the model’s ability to generalize to unseen data. Additionally, these mixing approaches help mitigate class imbalance by creating interpolated WSIs that better balance the representation of different classes.
Our ablation study demonstrates the robustness of our method to variations in slide mixing loss hyperparameters, showing that performance remains stable across different hyperparameter settings. This alleviates the need for extensive hyperparameter tuning, making the method practical and user-friendly. The choice of mixing locations within the MIL aggregator also plays a critical role in the effectiveness of these techniques. Specifically, incorporating Manifold Mixup within the Transformer Encoder layer provides the most significant performance improvement. This enhancement can be attributed to the richer and more abstract feature representations at this layer, where interpolations act as effective feature space regularization. By maintaining the contextual integrity of original WSIs, Manifold Mixup within the Transformer architecture ensures coherent learning and improved classification performance.
@@
We validate the proposed framework across various WSI training dataset sizes and labeling conditions. Under traditional fully-supervised fine-tuning with random sampling, our framework consistently outperforms the baseline. In addition, in an active learning setting, we evaluate its effectiveness using five acquisition functions: Entropy, K-means++, Coreset, BADGE, and CDAL. These functions are used to select the most informative WSIs for training which results in different dataset for training. Our results highlight that the proposed framework remains robust across varying WSI datasets and training sizes, demonstrating its versatility and effectiveness.
Limitations and Future Work. While our framework demonstrates significant performance improvements, the pre-training dataset is limited in terms of organ diversity and the number of WSIs. Expanding the dataset to include a broader range of organs and larger WSI collections would likely enhance the robustness and generalizability of the proposed methods. Despite these constraints, this study emphasizes the potential of the Barlow Twins Slide Mixing approach, showcasing the effectiveness of non-contrastive methods combined with intra-batch mixing for pre-training the MIL aggregator. By leveraging unlabeled WSIs, our framework provides a scalable solution for improving WSI classification, offering valuable insights for future research in weakly-supervised learning and histopathology.
@@ 6 Conclusions
MIL aggregator pre-training is a crucial yet underexplored area in weakly supervised WSI classification. Most existing approaches train MIL aggregators from scratch, overlooking the potential of unlabeled WSIs. This paper introduces PreMix, a framework designed to leverage extensive, underutilized unlabeled WSI datasets by employing Barlow Twins Slide Mixing for MIL aggregator pre-training. By incorporating Mixup and Manifold Mixup during fine-tuning, PreMix demonstrates robust performance across diverse WSI training datasets and labels, including traditional fully-supervised fine-tuning settings and five active learning techniques. The results highlight PreMix’s potential to enhance WSI classification with limited labeled data and its applicability to real-world histopathology practices.

### v3

@@ 3.4.1 Pre-training: Barlow Twins Slide Mixing
Initializing neural networks with pre-trained models offers the advantage of leveraging knowledge from unlabeled datasets, providing a strong starting point with weights that capture general patterns and representations. This is especially effective for tasks with limited labeled data [62].
@@
Slide augmentation. Standard augmentations from the PyTorch Torchvision library [63] are not directly applicable in our context, as our inputs are extracted WSI features rather than natural images. To address this limitation, we adapt four specific augmentations: random zeroing, Gaussian noise addition, random scaling, and random cropping, so they can be applied to WSI features during pre-training. As shown in Figure 3, for the $N^{\text{th}}$ WSI in the batch, the augmented features are denoted as $X_{N\text{A}}$ and $X_{N\text{B}}$. After applying augmentations to all WSIs in the batch, the features $X_{1\text{A}}$ to $X_{N\text{A}}$ are combined into $X_{\text{A}}$, and similarly, $X_{1\text{B}}$ to $X_{N\text{B}}$ are combined into $X_{\text{B}}$.
Intra-batch mixing. Inspired by prior work [59], which demonstrated the effectiveness of image mixing in MoCo V2 for improving downstream image classification, we introduce an intra-batch mixing strategy to the Barlow Twins method (a non-contrastive approach) to enhance the generation of positive slide pairs during MIL aggregator pre-training. Specifically, slide mixing is performed within a single batch to produce additional inputs, denoted as $X_{A-\text{Mix}}$ and $X_{B-\text{Mix}}$, as illustrated in Figure 3.
@@
The details of the OriginalLoss and MixingLoss functions, used to compute the Barlow Twins Slide Mixing loss, are provided in Algorithm 1. The OriginalLoss function calculates the standard Barlow Twins loss, while the MixingLoss function computes the loss between $Z_{B-{\text{Mix}}}$ and $Z_{A}$ or $Z_{B-{\text{Mix}}}$ and $Z_{A-{\text{Mix}}}$, accounting for $\lambda(i)$ as defined above. Here, “flip” refers to reversing the batch order.
@@
source_loss $\leftarrow\textsc{OriginalLoss}(Z_{A},Z_{B})$ (9)
The total loss is defined as follows:
@@ 4 Experiments
Table 1: WSI dataset distribution during pre-training and downstream classification phases. An “x” indicates that the dataset is not used in a particular process (e.g., pre-training, training, or testing), while a “-” indicates that distributions depend on the random sampling or active learning (AL) strategy used to select a 20 WSI sample budget per iteration from the unlabeled pool.
@@ 4.1 Data acquisition and preparation
This study utilizes the Camelyon16 dataset [64] [^1] Camelyon16 dataset, which contains 399 hematoxylin and eosin (H&E) stained whole slide images (WSIs) of breast cancer lymph node sections collected from two hospitals in the Netherlands: Radboud University Medical Center (RUMC) and University Medical Center Utrecht (UMCU). The dataset is divided into 270 training slides (159 normal and 111 tumor) and 129 test slides, serving as a widely used benchmark for WSI classification tasks.
For the pre-training phase, we supplement Camelyon16 with data from the Clinical Proteomic Tumor Analysis Consortium (CPTAC) Uterine Corpus Endometrial Carcinoma (UCEC) dataset [65] [^2] CPTAC UCEC dataset. From this dataset, 575 H&E stained WSIs of endometrial tissue samples are selected, comprising 185 normal and 390 tumor slides. The UCEC dataset focuses on endometrial cancer, a significant gynecological malignancy, providing diverse tissue samples for robust representation learning. During pre-training, both Camelyon16 and UCEC datasets are utilized in an unsupervised manner, with no access to slide-level labels to ensure a purely self-supervised learning process. All test data are excluded from pre-training to prevent information leakage.
To evaluate the model’s performance in data-limited scenarios, we restrict training to 100 WSIs from the Camelyon16 dataset. Initially, 20 WSIs are randomly selected for training, followed by iterative sample selection using either random sampling (traditional fully supervised fine-tuning) or active learning acquisition functions [36, 37, 38, 39, 40]. The complete Camelyon16 test set of 129 WSIs is retained for final evaluation. A detailed summary of the WSI dataset distribution across the pre-training and downstream classification phases is provided in Table 1.
@@ 4.2 Implementation setup
To operate under resource constraints and demonstrate the efficiency of our framework, all training experiments, including pre-training and downstream classification, are conducted on a single NVIDIA GeForce RTX 2080 Ti GPU. The models are implemented and trained using the PyTorch library [63], a widely used open-source deep learning platform, with all experiments performed in Python. For tiling WSIs, we use the OpenSlide library [66].
@@ 4.2.1 Region extraction
WSIs are tiled into regions of size $[4096,4096]$ pixels with no overlap between adjacent regions. A tissue threshold of 0.1 is applied to exclude non-informative regions. To standardize tissue patch extraction and address magnification differences across scanners in the Camelyon16 dataset, we employ HS2P [67], based on CLAM [17], with a pixel-spacing set to 0.5 μm/px (equivalent to 20$\times$ magnification).
@@ 4.3 Model evaluation
In the baseline framework, the MIL aggregator is initialized and trained from scratch. To evaluate the effectiveness of the proposed framework, we compute the average F1 for each active learning acquisition function and random sampling (traditional fully supervised fine-tuning) across label budgets. This comprehensive evaluation highlights the robustness and adaptability of our framework under varying dataset compositions and labeling budgets.
@@ 4.4 Impact of MIL aggregator pre-training methods
Building on the findings of [30], which employs the contrastive SimCLR method for MIL aggregator pre-training, we evaluate the non-contrastive Barlow Twins method alongside our proposed Barlow Twins Slide Mixing approach. The results, summarized in Table 2, reveal a clear difference in performance between these methods. Pre-training the MIL aggregator with SimCLR results in a 3.8% reduction in mean F1 compared to the baseline (training the aggregator from scratch), averaged across random sampling (traditional fully supervised fine-tuning) and five active learning acquisition functions.
In contrast, the vanilla Barlow Twins method yields a 1.4% improvement over the baseline, underscoring the potential of non-contrastive approaches for WSI pretraining. Building on this, our proposed Barlow Twins Slide Mixing enhances the original method by introducing intra-batch mixing, which creates additional positive WSI pairs during training. This simple yet effective modification leads to a further 2% gain in mean F1 score compared to the vanilla version, demonstrating the strength of our strategy for learning from unlabeled WSI data.
@@ 4.5 Impact of Mixup and Manifold Mixup integration
Table 3 presents the average performance across random sampling and five active learning acquisition functions under five WSI label budgets. Applying Mixup and Manifold Mixup during MIL aggregator fine-tuning, following Barlow Twins Slide Mixing pretraining, achieves the best results. This setup improves mean F1 by 1.3% over the no-mixing variant and 4.7% over the baseline. The gains are consistent across acquisition strategies and budgets, highlighting the robustness of our framework in low-label settings. These results demonstrate the importance of effective pretraining and the benefit of incorporating feature and label mixing during fine-tuning for improved WSI classification.
Table 3: Overall Results: Average performance of various acquisition functions under different budget constraints, evaluated with the F1 score. “SM” means Slide Mixing.
@@ 4.6.1 Effect of slide mixing loss hyperparameters
We investigate the impact of adjusting the hyperparameters $(\alpha,\beta,\gamma)$ in the Barlow Twins Slide Mixing method over 300 training epochs. These hyperparameters control the relative contributions of $source\_loss$, $mix\_source\_loss$, and $mix\_loss$ as defined in Equation 10. As shown in Table 4, the optimal configuration is $(1,0.5,0.5)$, which yields the highest mean F1 score. Other settings lead to slightly lower performance, indicating that Barlow Twins Slide Mixing maintains stable effectiveness even when the loss weights are moderately varied.
Table 4: Effect of adjusting hyperparameters $(\alpha,\beta,\gamma)$ in Barlow Twins Slide Mixing loss, corresponding to $source\_loss$, $mix\_source\_loss$, and $mix\_loss$ in Equation 10, respectively, evaluated with the F1 score.
@@ 4.6.2 Effect of Mixup and Manifold Mixup locations
We evaluate the impact of applying Mixup and Manifold Mixup at different stages of MIL aggregator fine-tuning. Specifically, we assess performance when mixing is applied at $Loc1$, $Loc2$, and $Loc3$, as shown in Figure 4. As reported in Table 5, applying mixing at all three locations yields the highest mean F1 score of 74.6%, outperforming all other configurations. In contrast, using only $Loc1$ leads to a slight performance drop, suggesting that mixing at a single stage is insufficient. Notably, applying Manifold Mixup at $Loc2$, within the Transformer Encoder, results in a 2.5% gain in mean F1, indicating that this layer plays a key role in improving classification accuracy.
@@ 5 Discussion
Most existing methods train the MIL aggregator from scratch, often overlooking the valuable information available in unlabeled WSIs. Our findings show that fine-tuning the MIL aggregator using contrastive learning methods such as SimCLR results in poorer performance compared to supervised training, which is consistent with previous studies [30]. The main drawback of contrastive learning lies in its reliance on both positive and negative pairs. In the context of WSI classification, where there is often a large imbalance between normal and tumor slides, this reliance increases the risk of including incorrect or noisy negative samples. To overcome this issue, we adopt non-contrastive methods such as Barlow Twins, which require only positive pairs. Unlike contrastive approaches, our experiments demonstrate that non-contrastive pre-training significantly improves WSI classification performance.
@@
We further extend the Barlow Twins method through our proposed Barlow Twins Slide Mixing, which introduces intra-batch mixing to create additional positive slide pairs during pre-training. By combining features from different slides within the same batch, this approach enhances semantic diversity and allows the MIL aggregator to learn more robust and generalizable representations.
To improve performance during fine-tuning, we apply Mixup and Manifold Mixup in a way that accounts for the unique properties of WSIs, including their large size and varying patch counts. These techniques interpolate both features and labels across slides, generating synthetic training examples that improve generalization. They also help address class imbalance by producing mixed slides that better represent underrepresented classes.
Our ablation study shows that the performance of our method remains stable across a range of hyperparameter values for slide mixing loss. This reduces the need for extensive hyperparameter tuning and makes the method more practical. In addition, we find that the location where mixing is applied within the MIL aggregator matters. In particular, applying Manifold Mixup within the Transformer Encoder layer yields the largest performance gain, likely due to the richer and more abstract representations learned at this stage. This location allows the model to benefit from interpolated features while preserving the contextual structure of the original slides.
@@
We validate our framework under different labeling conditions and training set sizes. Under traditional fully supervised fine-tuning with random sampling, our method consistently outperforms the baseline. We also evaluate it in an active learning setting using five acquisition strategies: Entropy, K-means++, Coreset, BADGE, and CDAL. These strategies select the most informative slides for training, resulting in different subsets of the data. Across all conditions, our method remains robust and effective, highlighting its adaptability and reliability in limited supervision settings.
Limitation and Future Work. While our method improves performance, the pre-training dataset is limited in organ diversity and slide count. In future work, we plan to expand the dataset to include a broader range of organs and cancer types to enhance generalizability.
@@ 6 Conclusions
MIL aggregator pre-training is a crucial yet underexplored area in weakly supervised WSI classification. Most existing approaches train MIL aggregators from scratch, overlooking the potential of unlabeled WSIs. This paper introduces PreMix, a framework designed to leverage extensive, underutilized unlabeled WSI datasets by employing Barlow Twins Slide Mixing for MIL aggregator pre-training. By incorporating Mixup and Manifold Mixup during fine-tuning, PreMix demonstrates robust performance across diverse WSI training datasets and labels, including traditional fully supervised fine-tuning settings and five active learning techniques. The results highlight PreMix’s potential to enhance WSI classification with limited labeled data and its applicability to real-world histopathology practices.